User Tools

Site Tools


handover:r1713

R1713

11:25 UT - Lost connection to all telescopes barring Ho and Hb. (Ellen)

Ke

  • Module NRAO-020, 0 GB
  • experiment did not start on time. First scan 313-2157a (Lucia)

Yg

  • Module USN-0173, 0 GB
  • experiment did not start on time. First scan 313-2157a (Lucia)
  • Scan 314-0301 had incorrect frequency settings. Some DBBC commands intended for Hb12 (under testing) were sent to Yg my mistake.
  • 07:51 UT - reported the error “ALARM: error m5 -104 mk5cn: time-out, connection closed”, which I've previously only seen for Ho 26m. As far as I can tell, again nothing is wrong. (Ellen)
  • 13:25 UT – VPN connection restored. Station looks to have operated OK in the break. (Dave)

Ho

  • experiment did not start on time. First scan 313-2156 (bad) (Lucia)
  • Receiver warm at the beginning, Brett turned the vacuum pump on. wrong receiver. First good scan 313-2258 on source 0748+126. (Lucia)
  • 1:02 UT: cnanot read receiver position correctly. But s/x is on. Receiver still warm, but on its way to get cooler. (Lucia)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1713.txt · Last modified: 2015/11/10 13:29 by Lucia Plank